I am working on my exercises and I am having problem with this question. I am very new to this course and having difficulty understanding these questions. Please kindly help me.
1. Find a trinomial that represents the area of a parallelogram whose base 3x +2 meters and whose height is 2x +3 meters.
2. Sam lives on a lot that he thought was a square, 157 feet. When he had it surveyed, he discovered that one side was actually 2 feet longer than he thought and the other was actually 2 feet shorter than he thought. How much less area does he have than he thought he had?
Thank you for your help.

Answer by checkley71(8403)

(3X+2)(2X+3)
--------------------------------
SQUARE LOT S^2=157 OR S=SQRT157 OR S=12.53 FEET FOR EACH SIDE.
(12.58-2)(12.58+2)=(10.58)(14.58)=154.26 SQUARE FEET.
THE DIFFERENCE IS 157-154.26=2.74 SQUARE FEET LOST.